OpenStudy (anonymous):

int sin(x^2) from 0 to inf i want full solution plz

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty} \sin (x^2) \ dx\] ??

OpenStudy (anonymous):

yes it is...

OpenStudy (anonymous):

just wait i'll write it for u on a paper

OpenStudy (turingtest):

can this even be done analytically?

OpenStudy (turingtest):

looks like it will have to with the Reimann-Zeta function

OpenStudy (anonymous):

yes i think just give me 2 min

OpenStudy (turingtest):

nice!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

tnx very nice solution

OpenStudy (anonymous):

yw :)